Could you help me with this problem?
Dina and Masha start out on a 10km bikepath at the same time. When Dina reaches the end of the 10km, Masha still has 2km left to bike. Dina's biking speed is 5km per hour, find Masha's biking speed.

Answer by checkley77(12844)

D=RT
10=5T
T=10/5
T=2 HOURS FOR DINA'S RIDE.
8=X*2
X=8/2
X=4 KPH IS MASKA'S SPEED.